# Multifractality of Many-Body Non-Hermitian Skin Effect

### v2

@@ I Noninteracting Hatano-Nelson model
II Interacting Hatano-Nelson model
@@
$H=\sum_{i=1}^{L}\left[2t\left(\left(1+\gamma\right)c_{i+1}^{{\dagger}}c_{i}+\left(1-\gamma\right)c_{i}^{{\dagger}}c_{i+1}\right)+4Jc_{i}^{{\dagger}}c_{i}c_{i+1}^{{\dagger}}c_{i+1}\right],$ (II.1)
@@
$H=\sum_{i=1}^{L}\left[\frac{t}{2}\left(\left(1+\gamma\right)\sigma_{i}^{-}\sigma_{i+1}^{+}+\left(1-\gamma\right)\sigma_{i}^{+}\sigma_{i+1}^{-}\right)+J\sigma_{i}^{z}\sigma_{i+1}^{z}\right],$ (II.2)
@@
which reduces to the non-Hermitian spin chain in the main text with no external fields (i.e., $g=h=0$). In contrast to the non-Hermitian spin chain in the main text with generic parameters, the interacting Hatano-Nelson model is invariant under $\mathrm{U}\left(1\right)$ spin rotation around the $z$ axis. In fact, it respects
$\left[H,\sum_{i=1}^{L}\sigma_{i}^{z}\right]=0.$ (II.3)
@@
Consequently, we study the multifractal scaling in a fixed subspace of the spin magnetization $\sum_{i=1}^{L}\sigma_{i}^{z}$. Below, we assume the even system length $L$ and focus on the half filling (i.e., $\sum_{i=1}^{L}\sigma_{i}^{z}=0$). It is also notable that this model should be integrable by the Bethe ansatz [112, 113, 114]. This contrasts with the non-Hermitian spin model introduced in the main text, which does not seem to be applicable to the Bethe ansatz and hence should be nonintegrable. The conformity of the spectral statistics to the random-matrix statistics also corroborates nonintegrability.
@@
We exactly diagonalize the interacting Hatano-Nelson model and obtain the complex spectrum, as well as the multifractal dimension for each right eigenstate [Fig. S1 (a, b)]. We choose the spin configuration as the computational basis of the participation entropy. To reduce numerical errors under the open boundary conditions, we first map Eq. (II.2) to the Hermitian XXZ chain through the similarity transformation [106, *Hatano-Nelson-97]. We then diagonalize the transformed Hermitian Hamiltonian and return to the original basis, obtaining the participation entropy of the interacting Hatano-Nelson model. In this procedure, no numerical instability arises since the diagonalization is concerned solely with Hermitian matrices. Under the periodic boundary conditions, the complex spectrum seems to be more structured than that of the nonintegrable non-Hermitian spin chain in the main text, which is also consistent with integrability of the interacting Hatano-Nelson model. Under the open boundary conditions, the entire many-body spectrum becomes real valued. Moreover, Fig. S1 (c, d) shows the distribution of multifractal dimensions as a function of the real part of many-body eigenenergies. In contrast to the nonintegrable model in the main text, multifractal dimensions no longer exhibit the characteristic behavior in which their peak clearly appears at the center of the many-body spectrum. Again, this should reflect integrability of the interacting Hatano-Nelson model. Furthermore, we provide the multifractal scaling in Fig. S2. In a similar manner to the nonintegrable model in the main text, the average participation entropy $\braket{S_{2}}$ decreases as non-Hermiticity $\gamma$ increases, showing the stronger many-body skin effect even in the presence of integrability. We also find that multifractal dimensions deviate from unity even under the periodic boundary conditions. This should also be a consequence of integrability, which leads to a departure from the random-matrix behavior.
Figure S1: Multifractality of the interacting Hatano-Nelson model ($t=1/\sqrt{2}$, $\gamma=0.6$, $J=1$). The particle number is chosen as $L/2$ (i.e., half filling). (a, b) Complex spectrum $\left(\mathrm{Re}\,E/L,\mathrm{Im}\,E/L\right)$ scaled by the system length $L=14$ under the (a) periodic boundary conditions (PBC) and (b) open boundary conditions (OBC). The color bars show the multifractal dimension $D_{2}=S_{2}/\ln\mathcal{N}$ for each right eigenstate. (c, d) Multifractal dimension $D_{2}$ as a function of $\mathrm{Re}\,E/L$ for the different system lengths $L$ under (c) PBC and (d) OBC.
Figure S2: Multifractality of the interacting Hatano-Nelson model ($t=1/\sqrt{2}$, $J=1$). The particle number is chosen as $L/2$ (i.e., half filling). (a, b) Participation entropy $\braket{S_{2}}$ averaged over all right eigenstates as a function of the Hilbert space dimension $\mathcal{N}$ under the (a) periodic boundary conditions (PBC) and (b) open boundary conditions (OBC) ($L=4,6,\cdots,14$). (c) Average multifractal dimensions $\braket{D_{2}}$ as functions of non-Hermiticity $\gamma$ under both PBC (red dots) and OBC (blue dots).
III Nonintegrable non-Hermitian spin chain
We provide additional numerical results for the nonintegrable non-Hermitian spin chain introduced in the main text:
$H=\sum_{i=1}^{L}\left[\frac{t}{2}\left(\left(1+\gamma\right)\sigma_{i}^{-}\sigma_{i+1}^{+}+\left(1-\gamma\right)\sigma_{i}^{+}\sigma_{i+1}^{-}\right)+J\sigma_{i}^{z}\sigma_{i+1}^{z}+g\sigma_{i}^{x}+h\sigma_{i}^{z}\right].$ (III.1)
To ensure the numerical precision, we perform the same calculations with the two different languages (C++ with LAPACK and Mathematica) and confirm the consistency between the two results. Additionally, we verify whether we can reproduce the original non-Hermitian Hamiltonian from the complex eigenvalues, along with the right and left eigenvectors, obtained from the numerical diagonalization.
In the absence of non-Hermiticity (i.e., $\gamma=0$), the eigenstates of Eq. (III.1) exhibit degeneracy, which is lifted upon the introduction of the non-Hermitian term ($\gamma\neq 0$). In such a case, any linear combination of the degenerate eigenstates remains an eigenstate, making the participation entropy explicitly dependent on the choice of the linear combination. Given that the degeneracy of eigenstates is lifted in the presence of non-Hermiticity, we here compute the participation entropy in the presence of infinitesimal non-Hermiticity $\gamma$ and regard this result as that for $\gamma=0$. We numerically confirm that the participation entropy remains essentially unchanged for $\gamma\lesssim 10^{-4}$ and the difference between $\gamma\to 0$ and small $\gamma\neq 0$ (e.g., $\gamma=0.2$) is negligible. It should also be noted that this procedure naturally selects eigenstates. For example, applying this method to the noninteracting Hatano-Nelson model without non-Hermiticity (i.e., $\gamma=0$) results in the selection of Bloch states, as the degeneracy for $\gamma=0$ is lifted.
III.1 Dependence on interaction strength
In Fig. S3, we provide additional numerical calculations of the multifractal dimensions for fixed non-Hermiticity $\gamma=0.8$ while changing the interaction strength $J$. Notably, the average multifractal dimensions $\braket{D_{2}}$ under the open boundary conditions remain nearly unchanged with some nonmonotonic oscillations even if we change $J$. This should imply that the non-Hermiticity $\gamma$ and many-body interaction $J$ compete with each other, leading to the formation of many-body skin modes.
Figure S3: Multifractality of the non-Hermitian spin chain in Eq. (III.1) ($t=1/\sqrt{2}$, $g=\left(5+\sqrt{5}\right)/8$, $h=\left(1+\sqrt{5}\right)/4$, $\gamma=0.8$). (a, b) Participation entropy $\braket{S_{2}}$ averaged over all right eigenstates as a function of the Hilbert space dimension $\mathcal{N}$ under the (a) periodic boundary conditions (PBC) and (b) open boundary conditions (OBC) ($L=8,9,\cdots,14$). (c) Average multifractal dimension $\braket{D_{2}}$ as functions of the interaction strength $J$ under both PBC (red dots) and OBC (blue dots).
III.2 Dependence on longitudinal field
In Fig. S4, we present additional numerical calculations for the subleading term $\braket{c_{q=2}}$ of the average participation entropy $\braket{S_{q=2}}$ for fixed non-Hermiticity $\gamma=0.1$ or $\gamma=0.6$ while changing the longitudinal field $h$. We numerically compute the participation entropy $\braket{S_{q=2}}$ averaged for all right many-body eigenstates and fit it by $\braket{S_{2}}=\braket{D_{2}}\ln\mathcal{N}+\braket{c_{2}}$. Under the periodic boundary conditions, the dependence of $\braket{c_{2}}$ on $h$ is similar for both small non-Hermiticity $\gamma=0.1$ and large non-Hermiticity $\gamma=0.6$. In contrast, under the open boundary conditions, $\braket{c_{2}}$ appears almost insensitive to $h$ for $\gamma=0.1$, whereas $\braket{c_{2}}$ exhibits a more pronounced change for $\gamma=0.6$. This significant difference may arise from the absence of the many-body skin effect for small non-Hermiticity, which merits further study.
Figure S4: Subleading term $\braket{c_{q=2}}$ of the average participation entropy $\braket{S_{q=2}}$ of the non-Hermitian spin chain in Eq. (III.1) ($t=1/\sqrt{2}$, $J=1$, $g=\left(5+\sqrt{5}\right)/8$) as a function of the longitudinal field $h$. The fitting by $\braket{S_{2}}=\braket{D_{2}}\ln\mathcal{N}+\braket{c_{2}}$ is performed up to $L=14$. We impose the periodic boundary conditions (PBC; red dots) or open boundary conditions (OBC; blue dots). The non-Hermitian term $\gamma$ is chosen to be (a) $\gamma=0.1$ or (b) $\gamma=0.6$.
III.3 $q$ dependence
In Fig. S5, we present the dependence of the multifractal dimension $\braket{D_{q}}$ on $q$. As shown in Fig. S5 (c), $\braket{D_{q}}$ exhibits the nontrivial $q$ dependence, which is a defining feature of multifractality.
Figure S5: Multifractality of the non-Hermitian spin chain in Eq. (III.1) ($t=1/\sqrt{2}$, $J=1$, $g=\left(5+\sqrt{5}\right)/8$, $h=\left(1+\sqrt{5}\right)/4$, $\gamma=0.4$). (a, b) $\braket{S_{q}}/\ln\mathcal{N}$ as a function of $1/\ln\mathcal{N}$ under the (a) periodic boundary conditions (PBC) and (b) open boundary conditions (OBC) ($L=8,9,\cdots,14$), where $\braket{S_{q}}$ denotes the $q$th participation entropy averaged over all right eigenstates, and $\mathcal{N}$ the Hilbert space dimension. (c) Average multifractal dimension $\braket{D_{q}}$ as functions of $q$ under both PBC (red dots) and OBC (blue dots).
III.4 Small non-Hermiticity
In Fig. S6, we provide additional numerical calculations of the multifractal dimensions for small non-Hermiticity $\gamma\lesssim 0.2$. Notably, the multifractal dimensions $D_{2}$ seem to be largely insensitive to the boundary conditions for $\gamma\lesssim 0.2$, which sharply contrasts with the strong boundary sensitivity for larger non-Hermiticity. This difference is consistent with the behavior of the average multifractal dimensions $\braket{D_{2}}$, as shown in the main text. The boundary insensitivity for small non-Hermiticity can arise from the absence of the non-Hermitian skin effect. Indeed, when the many-body interaction $J$ is stronger than non-Hermiticity $\gamma$, spins (or more precisely, quasiparticles of spin systems) are bounded by the many-body interaction and cannot exhibit the skin effect.
Figure S6: Multifractality of the non-Hermitian spin chain in Eq. (III.1) ($t=1/\sqrt{2}$, $J=1$, $g=\left(5+\sqrt{5}\right)/8$, $h=\left(1+\sqrt{5}\right)/4$, $L=14$). (a, b) Complex spectrum $\left(\mathrm{Re}\,E/L,\mathrm{Im}\,E/L\right)$ scaled by the system length $L=14$ under the (a) periodic boundary conditions (PBC) and (b) open boundary conditions (OBC) ($\gamma=0.8$). The color bars show the multifractal dimension $D_{2}=S_{2}/\ln\mathcal{N}$ for each right eigenstate. (c, d) Multifractal dimension $D_{2}$ as a function of $\mathrm{Re}\,E/L$ for different non-Hermiticity $\gamma$ under (c) PBC and (d) OBC.
III.5 Midspectrum eigenstates
In Fig. S7, we present additional numerical calculations of the multifractal dimensions averaged exclusively over the midspectrum right eigenstates satisfying $\left|\mathrm{Re}\,E/L\right|\leq 0.5$, instead of considering all the right eigenstates. Specifically, in Fig. S7 (e), we compare the multifractal dimensions averaged over all the eigenstates with those over the midspectrum eigenstates. The results show no significant deviations, thereby validating our use of the average multifractal dimensions to characterize the non-Hermitian skin effect.
Figure S7: Multifractality of the non-Hermitian spin chain in Eq. (III.1) ($t=1/\sqrt{2}$, $J=1$, $g=\left(5+\sqrt{5}\right)/8$, $h=\left(1+\sqrt{5}\right)/4$). (a, b) Participation entropy $\braket{S_{2}}$ averaged over the midspectrum eigenstates with $\left|\mathrm{Re}\,E/L\right|\leq 0.5$ as a function of the Hilbert space dimension $\mathcal{N}$ under the (a) periodic boundary conditions (PBC) and (b) open boundary conditions (OBC) ($L=8,9,\cdots,14$). (c, d) $\braket{S_{2}}/\ln\mathcal{N}$ as a function of $1/\ln\mathcal{N}$ under the (c) PBC and (d) OBC. (e) Average multifractal dimension $\braket{D_{2}}$ for all and midspectrum eigenstates as functions of non-Hermiticity $\gamma$ under the PBC and OBC.
IV Liouvillian skin effect
The many-body Lindbladian studied in the main text reads in the double Hilbert space
${\cal L}=-\text{i}\left({H}\otimes I-I\otimes{H}^{*}\right)+\sum_{i=1}^{L}\left[{L}_{i}\otimes{L}_{i}^{*}-\frac{1}{2}\,({L}_{i}^{{\dagger}}{L}_{i}\otimes I)-\frac{1}{2}\,(I\otimes{L}_{i}^{T}{L}_{i}^{*})\right],$ (IV.1)
where the Hamiltonian is
$H=\sum_{i=1}^{L}\left[\frac{t}{2}\left(\sigma_{i}^{-}\sigma_{i+1}^{+}+\sigma_{i}^{+}\sigma_{i+1}^{-}\right)+J\sigma_{i}^{z}\sigma_{i+1}^{z}+h\sigma_{i}^{z}\right],$ (IV.2)
and the dissipators are
$L_{i}=\sqrt{2\gamma}\sigma_{i}^{-}\sigma_{i+1}^{+}.$ (IV.3)
Here, we introduce participation entropy and multifractal dimensions based on the combination of right and left eigenoperators, instead of those based solely on right eigenoperators. We consider right and left eigenoperators of non-Hermitian superoperators,
|$\ket{r_{\alpha}}=\sum_{n=1}^{\cal N}r_{n}\ket{n},\quad\ket{l_{\alpha}}=\sum_{n=1}^{\cal N}l_{n}\ket{n}\quad\left(r_{n},l_{n}\in\mathbb{C}\right),$ (IV.4)
$\braket{r_{\alpha}}{r_{\alpha}}=\braket{l_{\alpha}}{l_{\alpha}}=1.$ (IV.5)
$S_{q}^{\rm RL}\coloneqq\frac{1}{1-q}\ln\left(\sum_{n=1}^{\cal N}\left|r_{n}l_{n}\right|^{q}\right),$ (IV.6)
and the multifractal dimension as
$D_{q}^{\rm RL}\coloneqq\frac{S_{q}^{\rm RL}}{\ln\mathcal{N}}=\frac{1}{1-q}\frac{1}{\ln\mathcal{N}}\ln\left(\sum_{n=1}^{\cal N}\left|r_{n}l_{n}\right|^{q}\right).$ (IV.7)
We calculate the participation entropy $\braket{S_{q=2}^{\rm RL}}$ averaged over all right and left eigenoperators (Fig. S8). As we increase the double Hilbert space dimension $\mathcal{N}$, the average participation entropy $\braket{S_{2}^{\rm RL}}$ under the periodic boundary conditions grows more slowly than $\braket{S_{2}^{\rm RL}}$ under the open boundary conditions. This behavior is qualitatively similar to the behavior of the average participation entropy $\braket{S_{2}}$ defined solely by right eigenoperators (see the main text). Additionally, as we increase the dissipation strength $\gamma$, the average participation entropy $\braket{S_{2}^{\rm RL}}$ remains almost the same for the periodic boundary conditions but increases for the open boundary conditions. It is worthwhile to further study the multifractal dimension $D_{q}^{\rm RL}$ defined by both right and left eigenoperators in a more systematic manner.
Figure S8: Multifractality of the Liouvillian skin effect ($t=1/\sqrt{2}$, $J=\left(5-\sqrt{3}\right)/3$, $h=\left(1+\sqrt{5}\right)/4$). In both bra and ket spaces, the particle number is chosen as $L/2$ and $\left(L-1\right)/2$ for even and odd $L$, respectively. Participation entropy $\braket{S_{2}^{\rm RL}}$ averaged over all right and left eigenoperators as a function of the double Hilbert space dimension $\mathcal{N}$ under the (a) periodic boundary conditions (PBC) and (b) open boundary conditions (OBC) ($L=5,6,\cdots,9$).

### v3

@@ Appendix A Noninteracting Hatano-Nelson model
Appendix B Nonintegrable non-Hermitian spin chain
We provide additional numerical results for the nonintegrable non-Hermitian spin chain in Eq. (5). To ensure the numerical precision, we perform the same calculations with the two different languages (C++ with LAPACK and Mathematica) and confirm the consistency between the two results. Additionally, we verify whether we can reproduce the original non-Hermitian Hamiltonian from the complex eigenvalues, along with the right and left eigenvectors, obtained from the numerical diagonalization.
In the absence of non-Hermiticity (i.e., $\gamma=0$), the eigenstates of Eq. (5) exhibit degeneracy, which is lifted upon the introduction of the non-Hermitian term ($\gamma\neq 0$). In such a case, any linear combination of the degenerate eigenstates remains an eigenstate, making the participation entropy explicitly dependent on the choice of the linear combination. Given that the degeneracy of eigenstates is lifted in the presence of non-Hermiticity, we here compute the participation entropy in the presence of infinitesimal non-Hermiticity $\gamma$ and regard this result as that for $\gamma=0$. We numerically confirm that the participation entropy remains essentially unchanged for $\gamma\lesssim 10^{-4}$ and the difference between $\gamma\to 0$ and small $\gamma\neq 0$ (e.g., $\gamma=0.2$) is negligible. It should also be noted that this procedure naturally selects eigenstates. For example, applying this method to the noninteracting Hatano-Nelson model without non-Hermiticity (i.e., $\gamma=0$) results in the selection of Bloch states, as the degeneracy for $\gamma=0$ is lifted.
Figure 5: Multifractality of the non-Hermitian spin chain in Eq. (5) ($t=1/\sqrt{2}$, $g=\left(5+\sqrt{5}\right)/8$, $h=\left(1+\sqrt{5}\right)/4$, $\gamma=0.8$). (a, b) Participation entropy $\braket{S_{2}}$ averaged over all right eigenstates as a function of the Hilbert space dimension $\mathcal{N}$ under the (a) periodic boundary conditions (PBC) and (b) open boundary conditions (OBC) ($L=8,9,\cdots,14$). (c) Average multifractal dimension $\braket{D_{2}}$ as functions of the interaction strength $J$ under both PBC (red dots) and OBC (blue dots).
B.1 Dependence on interaction strength
In Fig. 5, we provide additional numerical calculations of the multifractal dimensions for fixed non-Hermiticity $\gamma=0.8$ while changing the interaction strength $J$. Notably, the average multifractal dimensions $\braket{D_{2}}$ under the open boundary conditions remain nearly unchanged with some nonmonotonic oscillations even if we change $J$. This should imply that the non-Hermiticity $\gamma$ and many-body interaction $J$ compete with each other, leading to the formation of many-body skin modes.
Figure 6: Subleading term $\braket{c_{q=2}}$ of the average participation entropy $\braket{S_{q=2}}$ of the non-Hermitian spin chain in Eq. (5) ($t=1/\sqrt{2}$, $J=1$, $g=\left(5+\sqrt{5}\right)/8$) as a function of the longitudinal field $h$. The fitting by $\braket{S_{2}}=\braket{D_{2}}\ln\mathcal{N}+\braket{c_{2}}$ is performed up to $L=14$. We impose the periodic boundary conditions (PBC; red dots) or open boundary conditions (OBC; blue dots). The non-Hermitian term $\gamma$ is chosen to be (a) $\gamma=0.1$ or (b) $\gamma=0.6$.
B.2 Dependence on longitudinal field
In Fig. 6, we present additional numerical calculations for the subleading term $\braket{c_{q=2}}$ of the average participation entropy $\braket{S_{q=2}}$ for fixed non-Hermiticity $\gamma=0.1$ or $\gamma=0.6$ while changing the longitudinal field $h$. We numerically compute the participation entropy $\braket{S_{q=2}}$ averaged for all right many-body eigenstates and fit it by Eq. (6). Under the periodic boundary conditions, the dependence of $\braket{c_{2}}$ on $h$ is similar for both small non-Hermiticity $\gamma=0.1$ and large non-Hermiticity $\gamma=0.6$. In contrast, under the open boundary conditions, $\braket{c_{2}}$ appears almost insensitive to $h$ for $\gamma=0.1$, whereas $\braket{c_{2}}$ exhibits a more pronounced change for $\gamma=0.6$. This significant difference may arise from the absence of the many-body skin effect for small non-Hermiticity, which merits further study.
Figure 7: Multifractality of the non-Hermitian spin chain in Eq. (5) ($t=1/\sqrt{2}$, $J=1$, $g=\left(5+\sqrt{5}\right)/8$, $h=\left(1+\sqrt{5}\right)/4$, $\gamma=0.4$). (a, b) $\braket{S_{q}}/\ln\mathcal{N}$ as a function of $1/\ln\mathcal{N}$ under the (a) periodic boundary conditions (PBC) and (b) open boundary conditions (OBC) ($L=8,9,\cdots,14$), where $\braket{S_{q}}$ denotes the $q$th participation entropy averaged over all right eigenstates, and $\mathcal{N}$ the Hilbert space dimension. (c) Average multifractal dimension $\braket{D_{q}}$ as functions of $q$ under both PBC (red dots) and OBC (blue dots).
B.3 $q$ dependence
In Fig. 7, we present the dependence of the multifractal dimension $\braket{D_{q}}$ on $q$. As shown in Fig. 7 (c), $\braket{D_{q}}$ exhibits the nontrivial $q$ dependence, which is a defining feature of multifractality.
Figure 8: Multifractality of the non-Hermitian spin chain in Eq. (5) ($t=1/\sqrt{2}$, $J=1$, $g=\left(5+\sqrt{5}\right)/8$, $h=\left(1+\sqrt{5}\right)/4$, $L=14$). (a, b) Complex spectrum $\left(\mathrm{Re}\,E/L,\mathrm{Im}\,E/L\right)$ scaled by the system length $L=14$ under the (a) periodic boundary conditions (PBC) and (b) open boundary conditions (OBC) ($\gamma=0.8$). The color bars show the multifractal dimension $D_{2}=S_{2}/\ln\mathcal{N}$ for each right eigenstate. (c, d) Multifractal dimension $D_{2}$ as a function of $\mathrm{Re}\,E/L$ for different non-Hermiticity $\gamma$ under (c) PBC and (d) OBC.
B.4 Small non-Hermiticity
In Fig. 8, we provide additional numerical calculations of the multifractal dimensions for small non-Hermiticity $\gamma\lesssim 0.2$. Notably, the multifractal dimensions $D_{2}$ seem to be largely insensitive to the boundary conditions for $\gamma\lesssim 0.2$, which sharply contrasts with the strong boundary sensitivity for larger non-Hermiticity. This difference is consistent with the behavior of the average multifractal dimensions $\braket{D_{2}}$, as shown in Fig. 1 (g). The boundary insensitivity for small non-Hermiticity can arise from the absence of the non-Hermitian skin effect. Indeed, when the many-body interaction $J$ is stronger than non-Hermiticity $\gamma$, spins (or more precisely, quasiparticles of spin systems) are bounded by the many-body interaction and cannot exhibit the skin effect.
Figure 9: Multifractality of the non-Hermitian spin chain in Eq. (5) ($t=1/\sqrt{2}$, $J=1$, $g=\left(5+\sqrt{5}\right)/8$, $h=\left(1+\sqrt{5}\right)/4$). (a, b) Participation entropy $\braket{S_{2}}$ averaged over the midspectrum eigenstates with $\left|\mathrm{Re}\,E/L\right|\leq 0.5$ as a function of the Hilbert space dimension $\mathcal{N}$ under the (a) periodic boundary conditions (PBC) and (b) open boundary conditions (OBC) ($L=8,9,\cdots,14$). (c, d) $\braket{S_{2}}/\ln\mathcal{N}$ as a function of $1/\ln\mathcal{N}$ under the (c) PBC and (d) OBC. (e) Average multifractal dimension $\braket{D_{2}}$ for all and midspectrum eigenstates as functions of non-Hermiticity $\gamma$ under the PBC and OBC.
B.5 Midspectrum eigenstates
In Fig. 9, we present additional numerical calculations of the multifractal dimensions averaged exclusively over the midspectrum right eigenstates satisfying $\left|\mathrm{Re}\,E/L\right|\leq 0.5$, instead of considering all the right eigenstates. Specifically, in Fig. 9 (e), we compare the multifractal dimensions averaged over all the eigenstates with those over the midspectrum eigenstates. The results show no significant deviations, thereby validating our use of the average multifractal dimensions to characterize the non-Hermitian skin effect.
Appendix C Interacting Hatano-Nelson model
Figure 10: Multifractality of the interacting Hatano-Nelson model ($t=1/\sqrt{2}$, $\gamma=0.6$, $J=1$). The particle number is chosen as $L/2$ (i.e., half filling). (a, b) Complex spectrum $\left(\mathrm{Re}\,E/L,\mathrm{Im}\,E/L\right)$ scaled by the system length $L=14$ under the (a) periodic boundary conditions (PBC) and (b) open boundary conditions (OBC). The color bars show the multifractal dimension $D_{2}=S_{2}/\ln\mathcal{N}$ for each right eigenstate. (c, d) Multifractal dimension $D_{2}$ as a function of $\mathrm{Re}\,E/L$ for the different system lengths $L$ under (c) PBC and (d) OBC.
Figure 11: Multifractality of the interacting Hatano-Nelson model ($t=1/\sqrt{2}$, $J=1$). The particle number is chosen as $L/2$ (i.e., half filling). (a, b) Participation entropy $\braket{S_{2}}$ averaged over all right eigenstates as a function of the Hilbert space dimension $\mathcal{N}$ under the (a) periodic boundary conditions (PBC) and (b) open boundary conditions (OBC) ($L=4,6,\cdots,14$). (c) Average multifractal dimensions $\braket{D_{2}}$ as functions of non-Hermiticity $\gamma$ under both PBC (red dots) and OBC (blue dots).
@@ Appendix C Interacting Hatano-Nelson model
$H$ $=\sum_{i=1}^{L}\left[2t\left(\left(1+\gamma\right)c_{i+1}^{{\dagger}}c_{i}+\left(1-\gamma\right)c_{i}^{{\dagger}}c_{i+1}\right)\right.$
$\qquad\qquad\qquad\qquad\left.+4Jc_{i}^{{\dagger}}c_{i}c_{i+1}^{{\dagger}}c_{i+1}\right],$ (24)
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\left.+J\sigma_{i}^{z}\sigma_{i+1}^{z}\right],$ (25)
@@
which reduces to the non-Hermitian spin chain in Eq. (5) with no external fields (i.e., $g=h=0$). In contrast to Eq. (5) with generic parameters, the interacting Hatano-Nelson model is invariant under $\mathrm{U}\left(1\right)$ spin rotation around the $z$ axis. In fact, it respects
$\left[H,\sum_{i=1}^{L}\sigma_{i}^{z}\right]=0.$ (26)
@@
Consequently, we study the multifractal scaling in a fixed subspace of the spin magnetization $\sum_{i=1}^{L}\sigma_{i}^{z}$. Below, we assume the even system length $L$ and focus on the half filling (i.e., $\sum_{i=1}^{L}\sigma_{i}^{z}=0$). It is also notable that this model should be integrable by the Bethe ansatz [110, 111, 112]. This contrasts with Eq. (5), which does not seem to be applicable to the Bethe ansatz and hence should be nonintegrable. The conformity of the spectral statistics to the random-matrix statistics, shown in Fig. 2, also corroborates nonintegrability.
@@
We exactly diagonalize the interacting Hatano-Nelson model and obtain the complex spectrum, as well as the multifractal dimension for each right eigenstate [Fig. 10 (a, b)]. We choose the spin configuration as the computational basis of the participation entropy. To reduce numerical errors under the open boundary conditions, we first map Eq. (C) to the Hermitian XXZ chain through the similarity transformation [106, *Hatano-Nelson-97]. We then diagonalize the transformed Hermitian Hamiltonian and return to the original basis, obtaining the participation entropy of the interacting Hatano-Nelson model. In this procedure, no numerical instability arises since the diagonalization is concerned solely with Hermitian matrices. Under the periodic boundary conditions, the complex spectrum seems to be more structured than that of Eq. (5), which is also consistent with integrability of the interacting Hatano-Nelson model. Under the open boundary conditions, the entire many-body spectrum becomes real valued.
Moreover, Fig. 10 (c, d) shows the distribution of multifractal dimensions as a function of the real part of many-body eigenenergies. In contrast to Eq. (5), multifractal dimensions no longer exhibit the characteristic behavior in which their peak clearly appears at the center of the many-body spectrum. Again, this should reflect integrability of the interacting Hatano-Nelson model.
Furthermore, we provide the multifractal scaling in Fig. 11. In a similar manner to the nonintegrable model in Eq. (5), the average participation entropy $\braket{S_{2}}$ decreases as non-Hermiticity $\gamma$ increases, showing the stronger many-body skin effect even in the presence of integrability. We also find that multifractal dimensions deviate from unity even under the periodic boundary conditions. This should also be a consequence of integrability, which leads to a departure from the random-matrix behavior.